OpenStudy (anonymous):

Hey Can some one help me please? I'll give a Medal. What is the 50th term in this sequence that begins -6, 0, 6, 12

OpenStudy (anonymous):

@k_lynn

OpenStudy (mathmate):

to check, you can substitute one or two terms.

OpenStudy (anonymous):

That wasnt right i'll get you my answer chocies. 292 288 286 290

OpenStudy (anonymous):

@k_lynn

OpenStudy (mathmate):

Let's say the sequence is a1=-6 a2=0 a3=6 a4=12 We try to make a rule such that all four terms are satisfied. Common difference =12-6=6-0=....=6 so the sequence is an=6n+k, but we don't know k yet, so substitute the first term: a1=6(1)+k=-6 Solve for k to get k=-6-6(1)=-12 We have thus established that the rule is: \(\Large a_n=6n-12\) Check for n=4 a4=6(4)-12=24-12=12.. yay!

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

@kitty_eubanks

OpenStudy (mathmate):

For the 50th term, put n=50 in the rule! \(\Large a_{250} = ?\)

OpenStudy (anonymous):

so it would be 286?

OpenStudy (anonymous):

I think i'm doing something wrong?

OpenStudy (mathmate):

\(\Large a_{250} = 6*50-12 = ?\)